find x
when the long leg is 12, the short leg is x, and the hypotenuse is x+8

Answer by waynest(281)

c^2 = a^2 + b^2
c = x + 8, a = 12, b = x
(x + 8)^2 = 12^2 = x^2
x^2 + 16x + 64 = 144 + x^2
x^2 - x^2 + 16x + 64 -64 = 144 - 64 + x^2 - x^2
16x = 80
16x/16 = 80/16
x = 5
check:
(5 + 8)^2 = 12^2 + 5^2
13^2 = 169
169 = 169